OpenStudy (anonymous):

FRACTIONS PLEASE HELP! 1. 9/40 * (2/3) 2. -4/33 divided by -2/3

OpenStudy (anonymous):

When multiplying fractions, multiply tops with tops and bottoms with bottoms. Dividing fractions means multiply by the second fraction upside down (inverted.) You can reduce to lowest terms at the end or cross cancel before multiplying if that's easier for you.

OpenStudy (anonymous):

#2 as example: \[-\frac{4}{33} \div -\frac{2}{3} = -\frac{4}{33} \times -\frac{3}{2}\] \[ -\frac{4}{33} \times -\frac{3}{2} = \frac{-4 \times -3}{33 \times 2}\]